# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 2 Preliminaries
$\sqrt{1-\delta}\leq\|\mathbf{A}\mathbf{u}\|_{2}\leq\sqrt{1+\delta},\qquad\forall\mathbf{u}\in\mathcal{U}^{\rm(N)}=\mathcal{U}\cap\mathbb{S}^{n-1}.$
To be specific, we will focus on sparse recovery and the corresponding program of $\ell_{1}$-norm minimization (1.1). Therefore, we typically utilize the RIP over $\Sigma^{n}_{ts}$ for certain $t>0$ to imply the instance optimality. We shall work with RIP over $\Sigma^{n}_{2s}$ and utilize the following. (Note that RIP$(\Sigma^{n}_{2s},\delta)$ with $\delta<\frac{\sqrt{2}}{2}$ works, and we set $\delta=\frac{1}{3}$ just for concreteness.)
@@
[…]
@@
Gaussian Width & Metric Entropy: We need to work with two natural quantities that characterize the complexity of a set $\mathcal{K}$. The first one is the Gaussian width $\omega(\mathcal{K}):=\mathbbm{E}\sup_{\mathbf{u}\in\mathcal{K}}~{}\mathbf{g}^{\top}\mathbf{u}$, where $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$. The second one is the metric entropy $\mathscr{H}(\mathcal{K},r)=\log\mathscr{N}(\mathcal{K},r)$ where $\mathscr{N}(\mathcal{K},r)$ denotes the covering number of $\mathcal{K}$ under radius $r$, defined as the minimal number of radius-$r$ $\ell_{2}$-balls needed to cover $\mathcal{K}$. Metric entropy can be bounded in terms of the Gaussian width via Sudakov’s inequality [49, Coro. 7.4.3],
@@
for some absolute constant $C$. We also have Dudley’s inequality [49, Sec. 8.1] for the converse purpose. A notable difference is that the Gaussian width remains invariant after taking the convex hull, while the metric entropy under $o(1)$ covering radius could change significantly. For instance, the set $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n}$ (whose elements are known as the approximately $s$-sparse signals in $\mathbb{B}_{2}^{n}$) can be essentially viewed as the convex hull of $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ [41, Lem. 3.1]. Their Gaussian widths are of the same order [40, Sec. 2],
@@
The dependence on $r$ in (2.7) is tight in some regime; see [41, Sec. 3] and [44, Sec. 4.3.3]. In particular, the cardinality of an $r$-net for $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ logarithmically increases with $r^{-1}$, while that of $\sqrt{s}\mathbb{B}_{1}^{n}$ increases quadratically with $r^{-1}$.
@@
The following upper bound is a simple consequence of the matrix deviation inequality [49, Sec. 9.1] and will be of recurring use: if $\mathbf{A}$ has i.i.d. $\mathcal{N}(0,1)$ entries and $\mathcal{T}\subset\mathbb{R}^{n}$, then for any $t\geq 0$,
@@
Perturbation of Complex Phase: Under the convention $\frac{x}{0}=\infty$ for any $x\geq 0$, it holds for any $a,b\in\mathbb{C}$ that (e.g., [10, Lem. 8])
@@ 3 Main Results
We present our results for $\mathbf{x}^{\sharp}$ obtained by normalizing $\hat{\mathbf{x}}=\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\epsilon)$ in (1.8). Throughout the paper, we assume complex Gaussian $\bm{\Phi}$ with i.i.d. $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ entries (whose real part and imaginary part are independent) without always explicitly stating it. We will first study the performance of $\hat{\mathbf{x}}$ and then transfer this to $\mathbf{x}^{\sharp}$ via (2.12); thus, it is useful to identify the ground truth that $\hat{\mathbf{x}}$ approximates. This is a scaled version of $\mathbf{x}$ given by [10, 27]
@@
as it is easy to check $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$. (This means that $\mathbf{x}^{\star}$ is the point that satisfies the linear measurements in (1.8) in a noiseless case.) With $\kappa=\sqrt{\frac{\pi}{2}}$, $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ sharply concentrates about $1$ due to sub-Gaussian tail bounds, as we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(\frac{1}{\sqrt{m}})$ by (2.1). [^4] By (2.1), we have $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=\|\frac{1}{m}\sum_{i=1}^{m}(\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1)\|_{\psi_{2}}\leq\frac{C}{m}(\sum_{i=1}^{m}\|\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1\|_{\psi_{2}}^{2})^{1/2}\leq\frac{C_{1}}{\sqrt{m}}$. As a consequence, we have the sub-Gaussian tail bound $\mathbbm{P}(|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|\geq t)\leq 2\exp(-c_{2}mt)$ for any $t\geq 0$, hence $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough with probability at least $1-2\exp(-c_{3}m)$. We will use this observation in subsequent analysis. Hence, $\mathbf{x}^{\star}$ is in general very close to $\mathbf{x}$.
@@ 3.1 Instance Optimality
Our first result concerns the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ for arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ over a general cone $\mathcal{U}$.
@@
$m\geq C_{1}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right)$ (3.2)
@@
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U},C_{2}\eta\log^{1/2}(\eta^{-1})),\quad\forall\mathbf{x}\in\mathcal{K}$
holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
@@
Further, setting $\eta=\frac{c\delta}{\sqrt{\log(\delta^{-1})}}$ with small enough $c$, Theorem 3.1 yields the following: The matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy $\operatorname{RIP}(\Sigma^{n}_{2s},\delta)$ (w.h.p.) as long as $m=\Omega(\delta^{-2}\log^{2}(\delta^{-1})s\log(\frac{en}{s}))$. This improves on [10, Thm. 1], which requires $m=\Omega(\delta^{-4}s\log(\frac{n\log(mn)}{\delta s}))$ for the same purpose, in terms of log factors and the dependence on $\delta$. Indeed, the dependence on $\delta$ matches that of achieving RIP via a Gaussian matrix up to log factors (e.g., see [31]).
@@
More importantly, Theorem 3.1 applies to arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ with a number of measurements proportional to $\omega^{2}(\mathcal{K})$. To see this, by Sudakov’s inequality (2.4) and $\omega^{2}(\mathcal{K}_{(t)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})=4\omega^{2}(\mathcal{K})$ [49, Sec. 7.5.1], we find that (3.2) can be implied by the following based only on the Gaussian width,
$m\geq C_{1}^{\prime}\left(\frac{\omega^{2}(\mathcal{U}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K})}{\eta^{8}}\right)\quad\text{with large enough }C_{1}^{\prime}.$ (3.3)
@@
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3),\qquad\forall\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}.$
The distortion $1/3$ can be replaced by any given positive constant $\delta$, up to changes in the values of $C_{1},C_{2},c_{3}$.
The proof of Theorem 3.1 is based on covering and analogous to [10]. Nonetheless, [10] is restricted to $\mathcal{K}=\Sigma^{n,*}_{s}$ or at most other $\mathcal{K}$ with metric entropy logarithmically depending on the covering radius, and the techniques therein do not suffice for proving Corollary 1. We make a number of nontrivial modifications, with the most notable one being to introduce an additional index set when controlling the orthogonal term (see Appendix A.1). To preserve the presentation flow, we postpone the proof of Theorem 3.1 and the detailed discussions to Appendix A.
@@
We assume the event in Corollary 1 holds and consider any $\mathbf{x}\in\mathbb{S}^{n-1}$. If $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})$, then the event in Corollary 1 gives $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3)$, and further Proposition 2.1 implies
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 5\frac{\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}}=5\|\mathbf{x}^{\star}\|_{2}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
Finally, we use (2.12) to obtain
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},$
as claimed. If $\mathbf{x}\notin\mathbb{B}_{1}^{n}(\sqrt{2s})$, meaning that $\|\mathbf{x}\|_{1}>\sqrt{2s}$, then we let $\mathbf{x}_{[s]}=\mathrm{arg}\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{u}-\mathbf{x}\|_{2}$ and notice that $\|\mathbf{x}_{[s]}\|_{1}\leq\sqrt{s}$, and we have $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})=\|\mathbf{x}-\mathbf{x}_{[s]}\|_{1}\geq\|\mathbf{x}\|_{1}-\|\mathbf{x}_{[s]}\|_{1}\geq(\sqrt{2}-1)\sqrt{s}$. Therefore,
@@
While we focus on sparse recovery via basis pursuit, the generality of Theorem 3.1 in terms of $(\mathcal{K},\mathcal{U})$ allows for a straightforward generalization to other signal structures, such as PO-CS of low-rank matrices. Our subsequent technical results (Theorem 3.7, Lemmas 4.1, 4.2) are also presented in a similar manner. Also note that Theorem 3.2 directly generalizes to other instance optimal algorithms such as iterative hard thresholding (see [20, Sec. 6]).

### v3

@@ 2.1 Linear Compressed Sensing
$\sqrt{1-\delta}\leq\|\mathbf{A}\mathbf{u}\|_{2}\leq\sqrt{1+\delta},\qquad\forall\mathbf{u}\in\mathcal{U}^{\mathbb{S}}=\mathcal{U}\cap\mathbb{S}^{n-1}.$
[…]
@@
To recover block sparse signals $\mathbf{x}=(\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top})^{\top}\in\Sigma^{n_{1}}_{s_{1}}\times\Sigma^{n_{2}}_{s_{2}}\subset\mathbb{R}^{n}$ from $\mathbf{y}=\mathbf{A}\mathbf{x}+\bm{\epsilon}$, we can use a constrained weighted $\ell_{1}$ minimization,
@@
2.2 Gaussian Width & Metric Entropy
We need to work with two natural quantities that characterize the complexity of a set $\mathcal{K}$. The first one is the Gaussian width $\omega(\mathcal{K}):=\mathbbm{E}\sup_{\mathbf{u}\in\mathcal{K}}~\mathbf{g}^{\top}\mathbf{u}$, where $\mathbf{g}\sim\mathcal{N}(0,\mathbf{I}_{n})$. The second one is the metric entropy $\mathscr{H}(\mathcal{K},r)=\log\mathscr{N}(\mathcal{K},r)$ where $\mathscr{N}(\mathcal{K},r)$ denotes the covering number of $\mathcal{K}$ under radius $r$, defined as the minimal number of radius-$r$ $\ell_{2}$-balls needed to cover $\mathcal{K}$. Metric entropy can be bounded in terms of the Gaussian width via Sudakov’s inequality [50, Coro. 7.4.3],
@@ 2.2 Gaussian Width & Metric Entropy
for some absolute constant $C$. We also have Dudley’s inequality [50, Sec. 8.1] for the converse purpose. A notable difference is that the Gaussian width remains invariant after taking the convex hull, while the metric entropy under $o(1)$ covering radius could change significantly. For instance, the set $\sqrt{s}\mathbb{B}_{1}^{n}\cap\mathbb{B}_{2}^{n}$ (whose elements are known as the approximately $s$-sparse signals in $\mathbb{B}_{2}^{n}$) can be essentially viewed as the convex hull of $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ [43, Lem. 3.1]. Their Gaussian widths are of the same order [42, Sec. 2],
@@
The dependence on $r$ in (2.7) is tight in some regime; see [43, Sec. 3] and [41, Sec. 4.3.3]. In particular, the cardinality of an $r$-net for $\Sigma^{n}_{s}\cap\mathbb{B}_{2}^{n}$ logarithmically increases with $r^{-1}$, while that of $\sqrt{s}\mathbb{B}_{1}^{n}$ increases quadratically with $r^{-1}$.
2.3 Concentration Bound
@@
The following upper bound is a simple consequence of the matrix deviation inequality [50, Sec. 9.1] and will be of recurring use: if $\mathbf{A}$ has i.i.d. $\mathcal{N}(0,1)$ entries and $\mathcal{T}\subset\mathbb{R}^{n}$, then for any $t\geq 0$,
@@
2.4 Perturbation of Complex Phase
Under the convention $\frac{x}{0}=\infty$ for any $x\geq 0$, it holds for any $a,b\in\mathbb{C}$ that (e.g., [10, Lem. 8])
@@ 3 Main Results
We present our results for $\mathbf{x}^{\sharp}$ obtained by normalizing $\hat{\mathbf{x}}=\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8). Throughout the paper, we assume complex Gaussian $\bm{\Phi}$ with i.i.d. $\mathcal{N}(0,1)+\mathcal{N}(0,1)\mathbf{i}$ entries (whose real part and imaginary part are independent) without always explicitly stating it. We will first study the performance of $\hat{\mathbf{x}}$ and then transfer this to $\mathbf{x}^{\sharp}$ via (2.12); thus, it is useful to identify the ground truth that $\hat{\mathbf{x}}$ approximates. This is a scaled version of $\mathbf{x}$ given by [10, 27]
@@
as it is easy to check $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$. (This means that $\mathbf{x}^{\star}$ is the point that satisfies the linear measurements in (1.8) in a noiseless case.) With $\kappa=\sqrt{\frac{\pi}{2}}$, $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ sharply concentrates about $1$: by (2.1), we have
$\Big\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\Big\|_{\psi_{2}}=\Big\|\frac{1}{m}\sum_{i=1}^{m}(\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1)\Big\|_{\psi_{2}}\leq\frac{C}{m}\Big(\sum_{i=1}^{m}\|\kappa^{-1}|\bm{\Phi}_{i}^{*}\mathbf{x}|-1\|_{\psi_{2}}^{2}\Big)^{1/2}\leq\frac{C_{1}}{\sqrt{m}},$
which gives the sub-Gaussian tail bound
$\mathbbm{P}\Big(\big|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\big|\geq t\Big)\leq 2\exp(-c_{2}mt)\,,\quad\forall t\geq 0.$ (3.2)
Hence, $\mathbf{x}^{\star}$ is in general very close to $\mathbf{x}$. We will provide a uniform version of this fact in the subsequent Proposition 4.1.
Before proceeding, we provide a deterministic proof framework of our upper bounds. It is adapted from the original work [27].
Lemma 3.1 (Deterministic error bound).
Fix $\mathbf{x}\in\mathbb{S}^{n-1}$, $\bm{\Phi}\in\mathbb{C}^{m\times n}$ such that $\|\bm{\Phi}\mathbf{x}\|_{1}>0$ and suppose that we observe $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ that is a possibly noisy version of $\mathbf{z}:=\operatorname{sign}(\bm{\Phi}\mathbf{x})$. Suppose that we compute $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ with $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1},\varepsilon)$ in (1.8) with $\mathbf{A}_{\breve{\mathbf{z}}}$ defined in (1.7) and some $\varepsilon\geq 0$. If $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$ and
$\varepsilon\geq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big],$ (3.3)
then
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}+\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}.$
Note that the condition (3.3) can be relaxed to $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.
Proof.
By (1.7), $\mathbf{x}^{\star}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\mathbf{x}$ and $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, we have
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}\|_{2}=\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{x}\|_{2}$
$\leq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big]\leq\varepsilon.$
(In view of the above display, the condition (3.3) can be relaxed to $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$.) Combining with $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$, by Proposition 2.1 we obtain
$\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 7\varepsilon+5\frac{\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}}.$
By $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, $\mathbf{x}=\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}$, (2.12) and $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}$,
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ $=\Big\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\Big\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\max\{\|\hat{\mathbf{x}}\|_{2},\|\mathbf{x}^{\star}\|_{2}\}}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}$
$\leq\frac{14\varepsilon}{\|\mathbf{x}^{\star}\|_{2}}+\frac{10\sigma_{\ell_{1}}(\mathbf{x}^{\star},\Sigma^{n}_{s})}{\sqrt{s}\|\mathbf{x}^{\star}\|_{2}}=\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}+\frac{10\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}},$
as desired. ∎
@@ 3.1 Instance Optimality
For any $\mathbf{w}\in\mathbb{C}^{m}$, recall that the matrix $\mathbf{A}_{\mathbf{w}}$ is defined in (1.7). Our first result concerns the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathcal{K}\}$ for arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ over a general cone $\mathcal{U}$.
@@
$m\geq C_{1}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right)$ (3.4)
@@
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U},C_{2}\eta\log^{1/2}(\eta^{-1})),\quad\forall\mathbf{x}\in\mathcal{K},\quad\text{where }\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$
holds with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$.
@@
Further, setting $\eta=\frac{c\delta}{\sqrt{\log(\delta^{-1})}}$ with small enough $c$, Theorem 3.1 yields the following: The matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy $\operatorname{RIP}(\Sigma^{n}_{2s},\delta)$ (w.h.p.) as long as $m=\Omega(\delta^{-2}\log^{2}(\delta^{-1})s\log(\frac{en}{s}))$. This improves on [10, Thm. 1], which requires $m=\Omega(\delta^{-4}s\log(\frac{n\log(mn)}{\delta s}))$ for the same purpose, in terms of log factors and the dependence on $\delta$. Indeed, the dependence on $\delta$ matches that of achieving RIP via a Gaussian matrix up to log factors (e.g., see [32]).
@@
More importantly, Theorem 3.1 applies to arbitrary $\mathcal{K}\subset\mathbb{S}^{n-1}$ with a number of measurements proportional to $\omega^{2}(\mathcal{K})$. To see this, by Sudakov’s inequality (2.4) and $\omega^{2}(\mathcal{K}_{(t)})\leq\omega^{2}(\mathcal{K}-\mathcal{K})=4\omega^{2}(\mathcal{K})$ [50, Sec. 7.5.1], we find that (3.4) can be implied by the following based only on the Gaussian width,
$m\geq C_{1}^{\prime}\left(\frac{\omega^{2}(\mathcal{U}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K})}{\eta^{8}}\right)\quad\text{with large enough }C_{1}^{\prime}.$ (3.5)
@@
$\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3),\quad\forall\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1},$
where $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$. The distortion $1/3$ can be replaced by any given positive constant $\delta$, up to changes in the values of $C_{1},C_{2},c_{3}$.
The proof of Theorem 3.1 is based on covering techniques and is analogous to [10]. Nonetheless, [10] is restricted to $\mathcal{K}=\Sigma^{n,*}_{s}$ or at most other $\mathcal{K}$ with metric entropy logarithmically depending on the covering radius, and the techniques therein do not suffice for proving Corollary 1. We make a number of nontrivial modifications, with the most notable one being to introduce an additional index set when controlling the orthogonal term (see Appendix A.1). To preserve the presentation flow, we postpone the proof of Theorem 3.1 and the detailed discussions to Appendix A.
@@
We assume the event in Corollary 1 holds and consider any $\mathbf{x}\in\mathbb{S}^{n-1}$. If $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})$, then the event in Corollary 1 gives $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},1/3)$, and further Lemma 3.1 implies
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 10\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}.$
If $\mathbf{x}\notin\mathbb{B}_{1}^{n}(\sqrt{2s})$, meaning that $\|\mathbf{x}\|_{1}>\sqrt{2s}$, then we let $\mathbf{x}_{[s]}=\mathrm{arg}\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{u}-\mathbf{x}\|_{2}$ and notice that $\|\mathbf{x}_{[s]}\|_{1}\leq\sqrt{s}$, and we have $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})=\|\mathbf{x}-\mathbf{x}_{[s]}\|_{1}\geq\|\mathbf{x}\|_{1}-\|\mathbf{x}_{[s]}\|_{1}\geq(\sqrt{2}-1)\sqrt{s}$. Therefore,
@@
While we focus on sparse recovery via basis pursuit, the generality of Theorem 3.1 in terms of $(\mathcal{K},\mathcal{U})$ allows for a straightforward generalization to other signal structures, such as PO-CS of low-rank matrices. Our subsequent technical results (Theorem 3.7, Lemmas 4.2, 4.1) are also presented in a similar manner.